in how many ways can a committe cosisting of three mae and two women be chosen from seven men and five women
1
Expert's answer
2011-10-18T08:35:13-0400
There are C37=7!/(3!(7-3)!)=35 ways to choose 3 men from 7 men and C25=5!/(2!(5-2)!)=10 ways to choose 2 women from 5. So, the total number of ways is N=35+10=45